# Unmasking Trees for Tabular Data

### v2

@@ 1 Introduction
We address this shortfall by training gradient-boosted trees to autoregressively unmask features in random order, taking as inspiration the benefits of this training objective applied to tabular Transformer models (Gulati and Roysdon, 2024) (TabMT). This autoregressive approach, which we dub UnmaskingTrees, naturally performs conditional generation (i.e. imputation): at inference time, we simply fill in and condition on observed values, autoregressively generating the remaining missing values. This contrasts with tabular diffusion modeling, for which the RePaint inpainting algorithm (Lugmayr et al., 2022) is employed to mediocre effect (Jolicoeur-Martineau et al., 2024b). Because the predictor for a given feature must condition on varying subsets of the other features, the ability of gradient-boosted trees to handle missing features makes them a natural choice for autoregressive modeling. Hence, we maintain the tree-based approach of Jolicoeur-Martineau et al. (2024b), while replacing their tree-based regressors with our novel tree-based probabilistic predictors, which we turn to next.
While mean-estimating regression models are satisfactory for diffusion, for autoregression we must inject noise, and hence must estimate the entire conditional distribution of each feature. We therefore revisit the long-studied problem of (tabular) probabilistic prediction (Le et al., 2005; Meinshausen and Ridgeway, 2006). Because the conditional distribution is possibly multi-modal, parametric approaches such as XGBoostLSS (März, 2019), NGBoost (Duan et al., 2020), and PGBM (Sprangers et al., 2021) are poor choices for our setting. Meanwhile, quantization of a continuous variable can model its multi-modality, but at the cost of destroying either low-resolution or high-resolution information. A diffusion-based method, Treeffuser Beltran-Velez et al. (2024), was recently proposed to address these problems. However, as a diffusion method, it suffers from slow sampling and is unable to provide closed-form density estimates; furthermore, Treeffuser does not naturally model discrete outcomes. To address these problems, we propose BaltoBot, a balanced tree of boosted trees. For each individual variable, we recursively divide its output space with the kernel density integral (KDI) quantizer (McCarter, 2023) into a “meta-tree” of binary classifiers, which for us are gradient-boosted trees. This allows us to efficiently generate samples and estimate densities, because each sample follows only one path from root to leaf of the meta-tree. Performing regression with hierchical classification proved successful in computer vision object bounding box prediction (Li et al., 2020), but has been surprisingly underexplored in tabular ML.
@@
We showcase UnmaskingTrees on two tabular case studies, and on the benchmark of 27 tabular datasets presented by Jolicoeur-Martineau et al. (2024b). Most notably on this benchmark, our approach offers state-of-the-art performance on imputation and on generation given training data with missingness; and it has competitive performance on vanilla generation. We also demonstrate that BaltoBot is on its own a useful method for probabilistic prediction, showing its advantages on synthetic case studies. Finally, we provide open-source code with an easy-to-use sklearn-style interface at https://github.com/calvinmccarter/unmasking-trees. In addition to being a useful method for practitioners, we hope our work sparks conversations within the tabular ML community about whether diffusion is really all you need for tabular conditional generation.
@@ 2.1 UnmaskingTrees for tabular joint distribution modeling
UnmaskingTrees combines the gradient-boosted trees of ForestDiffusion (Jolicoeur-Martineau et al., 2024b) with the training objective of TabMT (Gulati and Roysdon, 2024), inheriting the benefits of both. Consider a dataset with $N$ examples and $D$ features. For each example, we generate new training samples by randomly sampling an order over the features, then incrementally masking the features in that random order. Given duplication factor $K$, we repeat this process times with $K$ different random permutations. This leads to a training dataset with $KND$ samples, given which we train XGBoost (Chen and Guestrin, 2016) models to predict each unmasked sample given the more-masked example derived from it. Implementing this is very simple: it requires about 70 lines of excessively-loquacious Python code for training, and about 20 lines for inference.
For both generation and imputation, we generate features of each sample in random order. For imputation rather than generation tasks, we begin by filling in each sample with the observed values, and run inference on the remaining unobserved features.
@@ 2.2 BaltoBot for tabular probabilistic prediction
Inspired by this, we propose a general-purpose solution to the tabular probabilistic prediction problem. For each individual regression output variable, we build a balanced tree of binary classifiers. Consider a node with depth $\delta$ on this “meta-tree”, which is fit with $(\mathbf{X}_{\textrm{train}}\in\mathbb{R}^{n\times d},\mathbf{y}_{\textrm{train}}\in\mathbb{R}^{n})$. Using kernel density integral quantization (KDI) (McCarter, 2023), which adaptively interpolates between uniform quantization and quantile quantization, we obtain binarized $\tilde{y}_{\textrm{train}}\in[0,1]^{n}$. We train an XGBoost classifier on $(\mathbf{X}_{\textrm{train}},\mathbf{\tilde{y}}_{\textrm{train}})$. If $\delta>0$, we then recursively pass $\{(\mathbf{x}^{(i)},y^{(i)})\in(\mathbf{X}_{\textrm{train}},\mathbf{y}_{\textrm{train}})|\tilde{y}^{(i)}=0\}$ to its left child, and analogously for $\tilde{y}^{(i)}=1$ to its right child. At a leaf node, $\delta=0$, if given a single unique training set output value in a bin, we record this value. At inference time, given a query input $\mathbf{x}$, we descend the tree by obtaining predicted probabilities from each node’s XGBoost classifier, then sampling from these. Once we reach a leaf node, we either sample uniformly from its appropriate bin, or we return the lone output value if a singleton bin.
At training and inference time, each XGBoost model within the meta-tree only sees examples that fall into its corresponding region of the output space. Thus, for a meta-tree with depth $\Delta$ (and thus $2^{\Delta}$ models), each example is only passed as input to $\Delta$ different models. While lower-level classifiers receive less data and are poorer quality, the magnitude of such errors are smaller due to our hierarchical partitioning approach. Furthermore, our singleton-bin technique allows us to adaptively generate discrete and even mixed-type variables, if these discrete outcomes are high-frequency relative to the total size of the data and to the depth of the meta-tree. (Up to $2^{\Delta}$ discrete outcomes can be produced by BaltoBot.) Finally, eschewing diffusion modeling enables us to perform closed-form conditional density estimation.
@@ 2.3 Computational complexity
ForestDiffusion, with $T$ diffusion steps and duplication factor $K$, constructs a training dataset of size $TKN\times D$. Given the same duplication factor $K$, UnmaskingTrees will construct a training dataset of size $KND\times D$. Meanwhile, ForestDiffusion must train $DT$ different XGBoost regression models. We, on the other hand, train $D$ different BaltoBot models, one per feature; with BaltoBot meta-tree depth of $\Delta$, we then train a total of $D2^{\Delta}$ XGBoost binary classifiers. However, classifiers lower in the BaltoBot meta-tree become progressively faster to train. Indeed, each constructed training sample will be seen by $DT$ different XGBoost regressors with ForestDiffusion, but only $D\Delta$ classifiers with our approach. Given that $T\sim 50$ and $\Delta\sim 4$, this yields a large speedup for our approach.
@@
At inference time, each ForestDiffusion generated sample passes through $T$ steps of the diffusion reverse-process, for a total of $DT$ XGBoost predictions. For UnmaskingTrees with BaltoBot, each generated sample instead requires only $D\Delta$ XGBoost predictions, because each sample follows only one path from root to leaf of the meta-tree. The resulting speedup is especially impactful for the multiple imputation scenario, where inference time dominates.
@@ 2.4 In-context learning-based generation with BaltoBoTabPFN and UnmaskingTabPFN
Within our flexible frameworks for joint and conditional modeling, TabPFN (Hollmann et al., 2022) can be used as a base learner for probabilistic prediction and generative modeling. For UnmaskingTabPFN joint modeling, a difficulty arises from TabPFN’s inability to handle inputs $\mathbf{X}_{\textrm{train}}$ with missing values. To address this, during both training and inference, we replaced NaNs with samples from $\mathcal{N}(0,1)$; we found this performed better than removing samples and/or features containing NaNs. To address TabPFN’s sample size limit of 1024, we performed random subsampling without replacement.
@@ 3 Results
Figure 2: Results on Iris dataset, with species, petal width, and petal length depicted. Original data and syntetically-generated datasets are shown on the left column. The imputed dataset is shown on the right column, with $\times$ symbols highlighting the samples with any missingness that required imputation.
We evaluate UnmaskingTrees on two case studies (Section 3.1) and on a tabular benchmark of 27 datasets (Section 3.2); we lastly evaluate BaltoBot on tabular probabilistic prediction (Section 3.3). Results were obtained always using the default hyperparameters: output tree depth of 4, and duplication factor $K=50$. Our hyperparameters were tuned on the two case study datasets, then applied without further tuning to the benchmark experiment. Overall, UnmaskingTrees has state-of-the-art performance on imputation and on generation after training on incomplete data; and it has competitive performance on vanilla tabular generation scenarios.
@@ 3.1 Case studies on Two Moons and Iris datasets
We first compare our approach to previous leading methods on the synthetic Two Moons dataset with 200 training samples and noise level $\mathcal{N}(0,0.1)$. We compare UnmaskingTrees to MissForest (Stekhoven and Bühlmann, 2012), MICE-Forest (Van Buuren et al., 1999; Wilson et al., 2022) (another leading traditional multiple imputation method), and ForestDiffusion, with default hyperparameters for all methods. For ForestDiffusion, we evaluate both the variance-preserving SDE (Forest-VP) and flow-matching (Forest-Flow) versions on generation; on imputation, we evaluate Forest-VP with and without RePaint, again using default RePaint hyperparameters.
We show results in Figure 1. On generation, Forest-VP appears to do best according to visual inspection, while UnmaskingTrees and Forest-Flow perform similarly decently. UnmaskingTabPFN performs poorly, but does capture the overall shape of the distribution. Next, we turn to imputation, wherein we request a single imputation for a copy of the original training data with the second dimension ($y$-axis) values masked out. ForestDiffusion struggles with and without RePaint, and MissForest and MICE-Forest have a lesser degree of out-of-distribution imputations. Meanwhile, UnmaskingTrees generates impeccable imputations.
@@ 3.2 Benchmarking UnmaskingTrees on 27 tabular datasets
We next repeat the experimental setup of Jolicoeur-Martineau et al. (2024b) for evaluating tabular imputation and generation methods. [^1] We do not compare against TabMT (Gulati and Roysdon, 2024) and TabPFGen (Ma et al., 2024) because no code was provided. Results for imputation are shown in Table 1. UnmaskingTrees wins first place on 3/9 metrics, including both metrics based on downstream prediction tasks; and it generally outperforms ForestDiffusion, winning on 8/9 metrics. While MissForest wins first place on 4/9 metrics, UnmaskingTrees wins 5-4 head-to-head against MissForest. UnmaskingTrees is also the only method with better than 5th place average ranking on all metrics. We report further ablation experiments in Appendix A, showing progressive improvements for the UnmaskingTrees framework, for KDI quantization versus k-Means, and for the BaltoBot method used in our full proposed solution.
We next repeat the experimental setup of Jolicoeur-Martineau et al. (2024b) for evaluating tabular generation methods. Results for partially-missing data are shown in Table 2. UnmaskingTrees is first place on 5/9 metrics; head-to-head, UnmaskingTrees beats TabDDPM 5-4, and beats Forest-Flow 6-3. Results for fully-observed data are shown in Table 3. UnmaskingTrees loses head-to-head to Forest-Flow, Forest-VP, and TabDDPM, but wins against the other methods.
@@
KNN 5.5 (0.5) 6.3 (0.4) 4.9 (0.4) 5 (0.4) 8.4 (0) 6.5 (1) 5.7 (1.1) 6.2 (1) 5.4 (0.6)
ICE 6.8 (0.4) 4.7 (0.4) 7 (0.5) 7.2 (0.4) 1.6 (0.2) 6.2 (1) 7 (0.6) 5.7 (0.9) 5.3 (0.6)
MICE-Forest 3.9 (0.4) 2.5 (0.4) 2.9 (0.2) 3 (0.2) 3.6 (0.2) 3.7 (1.4) 3.2 (1) 5.5 (1.2) 4.3 (0.6)
MissForest 2.7 (0.5) 4 (0.4) 1.8 (0.3) 2 (0.3) 5.5 (0.2) 3.8 (1.4) 2.5 (0.5) 5.5 (1.5) 3.3 (0.5)
Softimpute 6.7 (0.4) 7.6 (0.4) 7.1 (0.5) 7.3 (0.5) 8.4 (0) 6 (0.9) 7.8 (0.4) 6.3 (0.9) 6.7 (0.4)
OT 5.9 (0.4) 6.1 (0.3) 6 (0.5) 6 (0.5) 3.7 (0.3) 6.2 (0.5) 6.8 (0.6) 5.5 (0.8) 4.8 (0.5)
GAIN 4.7 (0.4) 6.5 (0.3) 6 (0.3) 6 (0.2) 6.9 (0.1) 5.7 (0.8) 5.4 (0.8) 4.7 (1) 5 (0.6)
Forest-VP 5.3 (0.4) 4 (0.5) 5.8 (0.3) 5.1 (0.4) 3.2 (0.4) 4.5 (0.9) 4.6 (0.8) 3.3 (0.6) 5.5 (0.7)
@@
Forest-Flow 3 (0.3) 2.6 (0.3) 2.6 (0.3) 2.7 (0.2) 3 (0.7) 3.7 (0.3) 5 (0.5) 3.8 (0.9) 3.2 (0.8)
UTrees 2.1 (0.2) 2.8 (0.3) 2.5 (0.2) 2.5 (0.2) 3.3 (0.8) 3.5 (0.5) 1 (0) 3.7 (0.9) 3.7 (1)
@@
UTrees 3.1 (0.1) 3.1 (0.2) 3.1 (0.2) 2.8 (0.2) 4.7 (0.3) 3.9 (0.3) 1 (0) 4.7 (0.7) 5.2 (0.9)

### v4

@@ 1 Introduction
We address this shortfall by training gradient-boosted trees to autoregressively unmask features in random order, via permutation language modeling (Yang, 2019). This autoregressive approach, which we dub UnmaskingTrees, naturally performs conditional generation (i.e. imputation): we simply fill in and condition on observed values, autoregressively generating the remaining missing values. This contrasts with tabular diffusion modeling, for which the RePaint inpainting algorithm (Lugmayr et al., 2022) is employed to mediocre effect (Jolicoeur-Martineau et al., 2024b). Because the predictor for a given feature must condition on varying subsets of the other features, the ability of gradient-boosted trees to handle missing features makes them a natural choice for autoregressive modeling. Hence, we maintain the tree-based approach of Jolicoeur-Martineau et al. (2024b), while replacing their tree-based regressors with our novel tree-based probabilistic predictors, which we turn to next.
While mean-estimating regression models are satisfactory for diffusion, for autoregression we must inject noise, and hence must estimate the entire conditional distribution of each feature. We therefore revisit the long-studied problem of (tabular) probabilistic prediction (Le et al., 2005; Meinshausen and Ridgeway, 2006). Because the conditional distribution is possibly multi-modal, parametric approaches such as XGBoostLSS (März, 2019), NGBoost (Duan et al., 2020), and PGBM (Sprangers et al., 2021) are poor choices for our setting. Meanwhile, quantization of a continuous variable can model its multi-modality, but at the cost of destroying either low-resolution or high-resolution information. A diffusion-based method, Treeffuser Beltran-Velez et al. (2024), was recently proposed to address these problems. However, as a diffusion method, it suffers from slow sampling and is unable to provide closed-form density estimates; furthermore, Treeffuser does not naturally model discrete outcomes. To address these problems, we propose BaltoBot, a balanced tree of boosted trees. For each individual variable, we recursively divide its output space with the kernel density integral (KDI) quantizer (McCarter, 2023) into a “meta-tree” of binary classifiers, which for us are gradient-boosted trees. This allows us to efficiently generate samples and estimate densities, because each sample follows only one path from root to leaf of the meta-tree. Performing regression with hierarchical classification proved successful in computer vision object bounding box prediction (Li et al., 2020), but has been surprisingly underexplored in tabular ML and in generative modeling.
@@
We showcase UnmaskingTrees on two tabular case studies, and on the benchmark of 27 tabular datasets presented by Jolicoeur-Martineau et al. (2024b). Most notably on this benchmark, our approach offers state-of-the-art performance on imputation and on generation given training data with missingness; and it has competitive performance on vanilla generation. We also demonstrate that BaltoBot is on its own a promising method for probabilistic prediction, showing its advantages on synthetic case studies and on a heavy-tailed sales forecasting benchmark.
Finally, we provide code with an easy-to-use sklearn-style API at https://github.com/calvinmccarter/unmasking-trees. In addition to being useful for practitioners, we hope our work sparks study within the tabular ML community about whether diffusion or autoregression is better for tabular data. Previous autoregressive tabular modeling methods, TabMT (Gulati and Roysdon, 2024) and DP-TBART (Castellon et al., 2023), use Transformer (Vaswani, 2017) models, making them less applicable for the GPU-poor. Our simple, efficient implementations of UnmaskingTrees and BaltoBot contribute to investigating this question.
@@ 2.1 UnmaskingTrees for tabular joint distribution modeling
UnmaskingTrees combines the gradient-boosted trees of ForestDiffusion (Jolicoeur-Martineau et al., 2024b) with the training objective of generalized autoregressive language modeling (Yang, 2019), inheriting the benefits of both. Consider a dataset with $N$ examples and $D$ features. We learn the joint distribution over $D$-dimensional example ${\mathbf{x}}$ by maximizing the expected log-likelihood with respect to all possible permutations of the factorization order,
$\log p({\mathbf{x}})=\log\mathbb{E}_{\sigma\in{\mathcal{U}}(G_{D})}\Big{[}\prod^{D}_{t=1}p\big{(}x_{\sigma(t)}|{\mathbf{x}}_{\sigma(<t)}\big{)}\Big{]},$
where $\sigma$ is a permutation drawn uniformly from ${\mathcal{U}}(G_{D})$, the permutation group on $D$ features; ${\mathbf{x}}_{\sigma(<t)}$ denotes all features that precede the $t$-th feature in the permuted sequence of features. If we were to have marginalized over permutations, we would have obtained a masked language modeling procedure with a randomly-sampled masking rate $r\sim{\mathcal{U}}(0,1)$ (Kitouni et al., 2023, 2024); such a procedure was previously shown to have benefits in combination with tabular Transformer models (Gulati and Roysdon, 2024) (TabMT).
For each example, we generate new training samples by randomly sampling an order over the features, then incrementally masking the features in that random order. Given duplication factor $K$, we repeat this process $K$ times with $K$ different random permutations, leading to a training dataset with $KND$ samples. Given this, we train XGBoost (Chen and Guestrin, 2016) models to predict each unmasked sample given the more-masked example derived from it, one per feature. We model categorical features via softmax-based classification with cross-entropy loss; our approach for continuous features is described in Section 2.2.
For both generation and imputation, we generate features of each sample in random order. For imputation rather than generation tasks, we begin by filling in each sample with the observed values, and run inference on the remaining unobserved features. Implementing this is very simple: it requires about 70 lines of Python code for training, and about 20 lines for inference.
@@ 2.2 BaltoBot for tabular probabilistic prediction
Inspired by this, we propose a general-purpose solution to the tabular probabilistic prediction problem. For each individual regression output variable, we build a height-$H$ balanced tree of binary classifiers. Consider a node with height $h$ on this “meta-tree”, which is fit with $({\mathbf{X}}_{\textrm{train}}\in{\mathbb{R}}^{n\times d},{\mathbf{y}}_{\textrm{train}}\in{\mathbb{R}}^{n})$. Using kernel density integral quantization (KDI) (McCarter, 2023), which adaptively interpolates between uniform quantization and quantile quantization, we obtain binarized $\tilde{{\mathbf{y}}}_{\textrm{train}}\in[0,1]^{n}$. Thus, the input space to every node is partitioned into two with the splitting point determined by KDI. We train an XGBoost classifier on $({\mathbf{X}}_{\textrm{train}},\mathbf{\tilde{y}}_{\textrm{train}})$. If $h>0$, we then recursively pass $\{({\mathbf{X}}^{(i)},y^{(i)})\in({\mathbf{X}}_{\textrm{train}},\mathbf{y}_{\textrm{train}})|\tilde{y}^{(i)}=0\}$ to its left child, and analogously for $\tilde{y}^{(i)}=1$ to its right child. At a leaf node, $h=0$, if given a single unique training set output value in a bin, we record this value. At inference time, given a query input ${\mathbf{X}}$, we descend the tree by obtaining predicted probabilities from each node’s XGBoost classifier, then sampling from these. Once we reach a leaf node, we either sample uniformly from its appropriate bin, or we return the lone output value if a singleton bin.
At training and inference time, each XGBoost model within the meta-tree only sees examples that fall into its corresponding region of the output space. Thus, for a meta-tree with height $H$ (and thus $2^{H}$ models), each example is only passed as input to $H$ different models. While lower-level classifiers receive less data and are poorer quality, the magnitude of such errors are smaller due to our hierarchical partitioning approach. Furthermore, our singleton-bin technique allows us to adaptively generate discrete and even mixed-type variables, if these discrete outcomes are high-frequency relative to the total size of the data and to the height of the meta-tree. (Up to $2^{H}$ discrete outcomes can be produced by BaltoBot.) Finally, eschewing diffusion modeling enables us to perform closed-form conditional density estimation.
Algorithms for UnmaskingTrees and BaltoBot are given in Appendix Section A.
@@ 2.3 Computational complexity
ForestDiffusion, with $T$ diffusion steps and duplication factor $K$, constructs a training dataset of size $TKN\times D$. Given the same duplication factor $K$, UnmaskingTrees will construct a training dataset of size $KND\times D$. Meanwhile, ForestDiffusion must train $DT$ different XGBoost regression models. We, on the other hand, train $D$ different BaltoBot models, one per feature; with BaltoBot meta-tree height of $H$, we then train a total of $D2^{H}$ XGBoost binary classifiers. However, classifiers lower in the BaltoBot meta-tree become progressively faster to train. Indeed, each constructed training sample will be seen by $DT$ different XGBoost regressors with ForestDiffusion, but only $DH$ classifiers with our approach. Given that $T\sim 50$ and $H\sim 4$, this yields a large speedup for our approach.
@@
At inference time, each ForestDiffusion generated sample passes through $T$ steps of the diffusion reverse-process, for a total of $DT$ XGBoost predictions. For UnmaskingTrees with BaltoBot, each generated sample instead requires only $DH$ XGBoost predictions, because each sample follows only one path from root to leaf of the meta-tree. The resulting speedup is especially impactful for the multiple imputation scenario, where inference time dominates.
@@ 2.4 In-context learning-based generation with BaltoBoTabPFN and UnmaskingTabPFN
Within our flexible frameworks for joint and conditional modeling, TabPFN (Hollmann et al., 2022) can be used as a base learner for probabilistic prediction and generative modeling. For UnmaskingTabPFN joint modeling, a difficulty arises from TabPFN’s inability to handle inputs ${\mathbf{X}}_{\textrm{train}}$ with missing values (NaNs). To address this, we developed NanTabPFN, a wrapper for TabPFN that supports missingness in both training and test features. Based on each test query ${\mathbf{x}}_{\textrm{test}}$, we select row indices ${\mathcal{R}}$ and column indices ${\mathcal{C}}$ so that $[{\mathbf{X}}_{\textrm{train}}]_{{\mathcal{R}},{\mathcal{C}}}$ has no NaNs, using the following key idea. Consider a particular train example ${\mathbf{x}}_{\textrm{train}}$ and test query ${\mathbf{x}}_{\textrm{test}}$, with visible (non-missing) features denoted by sets ${\mathcal{V}}({\mathbf{x}}_{\textrm{train}})$ and ${\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$. We can maximize the number of utilized features, while also ensuring that TabPFN receives no NaNs, by restricting the set of columns to those observed for the test query, ${\mathcal{C}}:={\mathcal{V}}({\mathbf{x}}_{\textrm{test}})$, then choosing training examples :=${\mathcal{R}}:=\{i|{\mathcal{C}}\subseteq{\mathcal{V}}({\mathbf{x}}^{(i)}_{\textrm{train}})\}$. In practice, our procedure is more complicated, because the above choices may result in either empty ${\mathcal{C}}$ or empty ${\mathcal{R}}$. If ${\mathcal{R}}$ is empty, we incrementally set random features of ${\mathbf{x}}_{\textrm{test}}$ to missing until we are able to obtain a non-empty training set. If ${\mathcal{C}}$ is empty, we introduce a new all-1s feature to both ${\mathbf{X}}_{\textrm{train}}$ and ${\mathbf{x}}_{\textrm{test}}$.
@@ 3 Results
Figure 2: Results on Iris dataset, with species, petal width, and petal length depicted. Original data and synthetically-generated datasets are shown on the left columns. The imputed dataset is shown on the right columns, with $\times$ symbols highlighting the samples with any missingness that required imputation.
We evaluate UnmaskingTrees on two case studies (Section 3.1) and on a tabular benchmark of 27 datasets (Section 3.2). We then evaluate BaltoBot and BaltoBoTabPFN on tabular probabilistic prediction case studies (Section 3.3) and on a sales forecasting dataset (Section 3.4). Results were obtained always using the default hyperparameters: output tree height of 4, and duplication factor $K=50$. These hyperparameter values were tuned on the Two Moons and Iris case studies, then applied without further tuning to the remaining experiments, because hyperparameter tuning is no fun at all. XGBoost hyperparameters were set to their defaults. Experiments were performed on a iMac (21.5-inch, Late 2015) with 2.8GHz Intel Core i5 processor and 16GB memory.
Overall, UnmaskingTrees (using BaltoBot) has state-of-the-art performance on imputation and on generation after training on incomplete data; and it has competitive performance on vanilla tabular generation scenarios. We further demonstrate the benefits of BaltoBot and BaltoBoTabPFN when evaluated in their own right for probabilistic prediction.
@@ 3.1 Case studies on Two Moons and Iris datasets
We first compare our approach to previous leading methods on the synthetic Two Moons dataset with 200 training samples and noise level $\mathcal{N}(0,0.1)$. We compare UnmaskingTrees to MissForest (Stekhoven and Bühlmann, 2012), MICE-Forest (Van Buuren et al., 1999; Wilson et al., 2022) (another popular traditional multiple imputation method), and ForestDiffusion, with default hyperparameters for all methods. For ForestDiffusion, we evaluate both the variance-preserving SDE diffusion (Forest-VP) and flow-matching (Forest-Flow) versions on generation; on imputation, we evaluate Forest-VP with and without RePaint, again using default RePaint hyperparameters; Forest-Flow does not support imputation.
We show results in Figure 1. On generation, Forest-VP appears to do best according to visual inspection, while UnmaskingTrees and Forest-Flow perform similarly decently. UnmaskingTabPFN performs poorly, but does capture the overall shape of the distribution. Next, we turn to imputation, wherein we request a single imputation for a copy of the original training data with the second dimension ($y$-axis) values masked out. ForestDiffusion struggles with and without RePaint, with substantial out-of-distribution imputations, and MissForest and MICE-Forest share this problem to lesser degrees. Meanwhile, UnmaskingTrees generates impeccable imputations.
@@ 3.2 Benchmarking UnmaskingTrees on 27 tabular datasets
Imputation
Here, we add UnmaskingTrees to the benchmark of 8 imputation methods on 27 public datasets, evaluated according to 9 metrics, developed by Jolicoeur-Martineau et al. (2024b) for evaluating tabular imputation and generation methods. This benchmark primarily contains smaller-sized (with $640103\leq N\leq 20{,}$ 640 and $4\leq D\leq 90$) datasets, which our approach is especially geared towards. Namely, we compare our approach against Forest-VP Jolicoeur-Martineau et al. (2024b), as well as k-NN imputation (Troyanskaya et al., 2001), ICE (Buck, 1960), MICE-Forest (Van Buuren et al., 1999; Wilson et al., 2022), MissForest (Stekhoven and Bühlmann, 2012), Softimpute (Hastie et al., 2015), minibatch Sinkhorn optimal transport (Muzellec et al., 2020), and generative adversarial nets (GAIN) (Yoon et al., 2018).  [^1] We do not add TabMT (Gulati and Roysdon, 2024) and TabPFGen (Ma et al., 2024) to the benchmark because no code was provided. We do not add UnmaskingTabPFN because of out-of-memory errors on our machine. We follow Jolicoeur-Martineau et al. (2024b) in computing the per-dataset rank of each method relative to other methods, then reporting the average over 27 datasets. For all methods other than our own, we compute ranks by reusing the raw scores provided in Jolicoeur-Martineau et al. (2024b)’s code repository.
Results for imputation are shown in Table 1. UnmaskingTrees wins first place on 3/9 metrics, including both metrics based on downstream prediction tasks; and it generally outperforms ForestDiffusion, winning on 8/9 metrics. While MissForest wins first place on 4/9 metrics, UnmaskingTrees wins 5-4 head-to-head vs MissForest; UnmaskingTrees has average averaged rank of 3.2 compared to 3.5 for MissForest. UnmaskingTrees is also the only method with better than 5th place rank on all metrics.
We report further ablation experiments in Table 2, wherein we run UnmaskingTrees without BaltoBot, and instead with vanilla quantization using k-Means clustering (Lloyd, 1982) and KDI quantization (McCarter, 2023). Results showing progressive improvements for the UnmaskingTrees framework, for KDI quantization versus k-Means, and for the BaltoBot method used in our full proposed solution.
@@
[…]
KNN 5.5 (0.5) 6.3 (0.4) 4.9 (0.4) 5.0 (0.4) 8.4 (0) 6.5 (1) 5.7 (1.1) 6.2 (1) 5.4 (0.6)
ICE 6.8 (0.4) 4.7 (0.4) 7.0 (0.5) 7.2 (0.4) 1.6 (0.2) 6.2 (1) 7.0 (0.6) 5.7 (0.9) 5.3 (0.6)
MICE-Forest 3.9 (0.4) 2.5 (0.4) 2.9 (0.2) 3.0 (0.2) 3.6 (0.2) 3.7 (1.4) 3.2 (1) 5.5 (1.2) 4.3 (0.6)
MissForest 2.7 (0.5) 4.0 (0.4) 1.8 (0.3) 2.0 (0.3) 5.5 (0.2) 3.8 (1.4) 2.5 (0.5) 5.5 (1.5) 3.3 (0.5)
Softimpute 6.7 (0.4) 7.6 (0.4) 7.1 (0.5) 7.3 (0.5) 8.4 (0) 6.0 (0.9) 7.8 (0.4) 6.3 (0.9) 6.7 (0.4)
OT 5.9 (0.4) 6.1 (0.3) 6.0 (0.5) 6.0 (0.5) 3.7 (0.3) 6.2 (0.5) 6.8 (0.6) 5.5 (0.8) 4.8 (0.5)
GAIN 4.7 (0.4) 6.5 (0.3) 6.0 (0.3) 6.0 (0.2) 6.9 (0.1) 5.7 (0.8) 5.4 (0.8) 4.7 (1) 5.0 (0.6)
Forest-VP 5.3 (0.4) 4.0 (0.5) 5.8 (0.3) 5.1 (0.4) 3.2 (0.4) 4.5 (0.9) 4.6 (0.8) 3.3 (0.6) 5.5 (0.7)
@@
KNN 6.8 (0.6) 7.8 (0.6) 6.0 (0.4) 6.1 (0.5) 10.4 (0) 8.2 (1.3) 7.0 (1.5) 7.5 (1.5) 6.5 (0.8)
ICE 8.3 (0.5) 5.8 (0.5) 8.5 (0.6) 8.8 (0.5) 1.9 (0.4) 8.0 (1.1) 9.0 (0.6) 7.2 (1.1) 6.4 (0.8)
MissForest 3.3 (0.7) 5.0 (0.6) 2.2 (0.4) 2.3 (0.4) 7.2 (0.3) 4.7 (1.8) 3.3 (0.9) 6.8 (1.9) 3.8 (0.6)
Forest-VP 6.4 (0.5) 4.8 (0.6) 7.0 (0.4) 6.1 (0.5) 3.8 (0.5) 6.5 (0.9) 6.6 (0.8) 4.5 (0.8) 6.5 (0.8)
UTrees-kMeans 6.0 (0.6) 5.8 (0.5) 6.3 (0.6) 6.1 (0.6) 4.1 (0.3) 4.0 (0.7) 2.9 (0.6) 3.8 (1) 6.0 (0.7)
UTrees-KDI 5.1 (0.5) 5.1 (0.5) 5.4 (0.6) 5.6 (0.5) 4.8 (0.3) 4.5 (0.9) 4.0 (0.5) 3.5 (1.2) 6.4 (0.7)
UTrees 3.8 (0.5) 3.2 (0.5) 3.8 (0.4) 3.8 (0.5) 5.0 (0.3) 2.7 (0.6) 2.9 (0.8) 3.5 (0.8) 5.8 (0.7)
Generation with and without missingness
We next repeat the experimental setup of Jolicoeur-Martineau et al. (2024b) for evaluating tabular generation methods. For tabular generation, using the same 27 datasets, Jolicoeur-Martineau et al. (2024b) benchmark their methods (Forest-VP and Forest-Flow) against 6 other methods, namely, Gaussian Copula (Joe, 2014), tabular variational autoencoding (TVAE) (Xu et al., 2019), two conditional generative adversarial net methods (CTGAN (Xu et al., 2019) and CTAB-GAN+ (Zhao et al., 2021)), and two other tabular diffusion methods (STaSy (Kim et al., 2022) and TabDDPM (Kotelnikov et al., 2023)). These are evaluated with 9 metrics, in the vanilla fully-observed setting and in the synthetically-induced 20% missing completely at random (MCAR) setting.
Results for partially-missing data are shown in Table 3. UnmaskingTrees is first place on 5/9 metrics; head-to-head, UnmaskingTrees beats TabDDPM 5-4, and beats Forest-Flow 6-3. Results for fully-observed data are shown in Table 4. UnmaskingTrees loses head-to-head to Forest-Flow, Forest-VP, and TabDDPM, but wins against the other methods.
@@
Forest-Flow 3.0 (0.3) 2.6 (0.3) 2.6 (0.3) 2.7 (0.2) 3.0 (0.7) 3.7 (0.3) 5.0 (0.5) 3.8 (0.9) 3.2 (0.8)
UTrees 2.1 (0.2) 2.8 (0.3) 2.5 (0.2) 2.5 (0.2) 3.3 (0.8) 3.5 (0.5) 1.0 (0.0) 3.7 (0.9) 3.7 (1.0)
@@
UTrees 3.1 (0.1) 3.1 (0.2) 3.1 (0.2) 2.8 (0.2) 4.7 (0.3) 3.9 (0.3) 1.0 (0.0) 4.7 (0.7) 5.2 (0.9)
Raw scores, per-dataset results, and runtimes are provided in the Appendix.